OpenStudy (anonymous):

You roll 2 dice. What is the probability that the sum of the dice is greater than 7 and 1 die shows a 2? A 6 X 6 table of dice outcomes will help you to answer this question.

OpenStudy (anonymous):

sum is greater than 7 AND one die shows a two?

OpenStudy (anonymous):

we can list them and count (2,6) or (6,2) that is all of them

OpenStudy (anonymous):

and as the 6 by 6 table shows there are \(6\times 6=36\) possibilities total, so your answer is \(\frac{2}{36}=\frac{1}{18}\)